# Ultra-low-power ring-based wireless mouse

### v1

@@ 1. INTRODUCTION
Mouse devices are basic tools for simple, fast pointing with computers (Villar et al., 2009). Integrating mouse functionality into wearable form factors such as wristbands, glasses and rings (e.g., Apple Watch, Galaxy Ring) promises always-available essential interactions with wearable computing platforms (Xu et al., 2023; Streli et al., 2024; Kienzle et al., 2021). Especially, the wireless ring-formed input devices worn on the index finger offer reliable detection of even subtle, privacy-preserved thumb-to-index finger inputs, unlike the wristband and glasses (Takahashi et al., 2024; Ogata et al., 2012; Kim et al., 2024). However, the physical constraint of the tiny ring structure requires the equipment of tens of small battery, challenging the continuous operation of tens of -class power-consuming wireless communication modules. For example, prior wireless rings support below ten hours of continuous wireless communication with Bluetooth Low Energy (BLE), limiting the ring’s communication usage to intermittent data transfer (Zhou et al., 2023; Kim et al., 2024; Shen et al., 2024; Jing et al., 2012). Such an operation is suitable for periodic healthcare monitoring around the rings (e.g., Oura Ring, Galaxy Ring), but not appropriate for ring-based input interface that requires long-term, real-time communication with other wearables (e.g., smartwatch, HMD).
This paper presents picoRing mouse, enabling an ultra-low-power ring-shaped wireless mouse with a ring-to-wristband coil-based impedance sensing (see Figure 1). picoRing mouse is inspired by a coil-based sensitive impedance sensing named passive inductive telemetry (PIT) (Takahashi et al., 2024, 2020). Unlike long-range electromagnetic communication such as BLE and RF backscatter, PIT constructs a short-range but ultra-low-powered inductive link between a pair of a ring coil and a wristband coil. Since the ring coil can send its sensor information to the wristband coil by simply modifying the inductive field generated from the wristband coil, the ring coil does not need active signal transmission with power-hungry communication modules. With the combination of PIT with the ring-based mouse module, picoRing mouse enables the ultra-low-powered mouse ring with maximum power consumption of approximately , enabling continuous operation of approximately (24hrs use/day)- (4hrs use/day) days on a single charge of a curved Lipo battery. To facilitate replication and provide comprehensive implementation details, we have open-sourced current picoRing mouse: https://removed-for-double-blind-review
@@ 2.2. Wireless Ring
Ring-formed mouse devices have long been investigated within HCI community. The exploration of the previous mouse ring can be divided into two categories: i) ring-based sensing techniques such as computer vision (Sun et al., 2021; Chan et al., 2015; Kienzle and Hinckley, 2014), inertial sensors (Shen et al., 2024; Liang et al., 2021), ultrasound (Yu et al., 2024; Zhang et al., 2017), and electromagnetic wave (Chen et al., 2022; Waghmare et al., 2023) and ii) low-powered wireless ring design using RF backscatter (Liu et al., 2013; Ensworth and Reynolds, 2017; Bainbridge and Paradiso, 2011), near-field communication (NFC) (Zhan et al., 2024; Lee et al., 2022; Liang and Guo, 2021), intrabody communication (Fukumoto and Tonomura, 1997; Varga et al., 2018), and magnet-based tracking (Ashbrook et al., 2011). The ring-based finger sensing demonstrates high-fidelity microgesture recognition compared to the wristband- and glasses-based finger sensing (Kienzle et al., 2021). However, the long-term operation of the low-powered wireless rings is still challenging (Wagih et al., 2023). Specifically, -class RF backscatter and magnet-mounted ring need large antenna and bulky magnet, which is hard to fit in the small ring (Naderiparizi et al., 2018; Ashbrook et al., 2011). Moreover, the magnet-based tracking via exposed static magnetic field is easily influenced by the surrounding magnetic noises such as electrical appliances, impairing the tracking accuracy of the magnet position or orientation (Wang et al., 2022). -class NFC supports small tag design, but the communication range is a few centimeter-scale (Gummeson et al., 2014; Takahashi et al., 2025; Zhang et al., 2022); such a short-range link cannot construct ring-to-wristband communication. Intrabody communication (Fukumoto and Tonomura, 1997; Varga et al., 2018), which uses capacitive body as a signal path, can transmit signal with -class ultra-low-power throughout the body (Li et al., 2021), but both the RF presence of other electronic devices and the daily items such as wooden desk and wall can impair the capacitive signal path, causing signal degradation, reduced transmission range, or even data loss (Kienzle et al., 2021).
@@ 3. SYSTEM DESIGN
picoRing mouse consists of two main components (see Figure 3a): 1) a ring coil with a mouse module, which changes its frequency response based on the thumb-to-index inputs, and 2) a wristband coil that detects the unique peak in the frequency response generated by the ring’s resonance. The working principle of picoRing mouse is as follows: First, the wristband coil generates a weak inductive field to couple with the nearby ring coil. When the user scrolls or presses the mouse module mounted on the ring, the MCU connected to the mouse module detects either the scrolling or pressing via the magnet or tactile switch, transforming the mouse input into a unique frequency response (i.e., the change in the ring’s resonant frequency) via a voltage-controlled variable capacitor (see Figure 3b). Owing to the inductive coupling, the wristband coil connected to a sensitive impedance measurement module obtains the frequency response. Since the ring’s frequency response is appeared as a sharp frequency peak at the ring’s resonant frequency in the frequency characteristics of wristband’s impedance, the wristband coil enables to recognize the mouse input (i.e., pressing and 2D scrolling) as a unique peak output.
@@ 3.1. Ring-to-wristband Wireless Communication
picoRing mouse extends PIT system in semi-passive manner. While the prior PIT features the full-passive design of the ring coil by modifying the ring-to-wristband inductive field with a full-passive modulation circuit composed of tactile switches. In contrast, picoRing mouse employs semi-passive modulation approach, which modifies the inductive field with a low-powered digital modulation circuit composed of the battery-assisted voltage-controlled variable capacitor. Note that semi-passive modulation approach does not use the active signal transmission circuit. Unlike full-passive modulation, the semi-passive modulation can transmits the sensor information as digital signals. Specifically, picoRing mouse uses a simple frequency-shift keying, which encodes the multi-modal mouse input into a corresponding frequency shift based on the voltage-controlled variable capacitor.
Let us explain how the ring coil can influence the wristband coil in the frequency-shifting manner. When the ring coil resonant at inductively couples with the wristband coil, the input impedance of the wristband coil, , can be expressed as follows:
where and the total impedance of the wristband and ring coils, respectively, the angular frequency, the mutual inductance between the ring and wristband coils. can be simplified as at because is in the resonance state. As a result, Equation 2 can be also simplified at as follows:
Equation 5 indicates increases significantly around , resulting in the sharp peak in the frequency characteristics of around . However, similar to picoRing (Takahashi et al., 2024), the extremely weak inductive coupling makes too small to detect through standard impedance measurement circuit.
@@
where is the input voltage of the bridge circuit, is the amplifier factor. Note that is valid under condition because has the large imaginary part at non-resonant frequency and is so small in the situation where the ring coil is away from the wristband coil. Equation 8 indicates that the wristband coil can reliably detect the ring’s resonant frequency through output () of the bridge circuit. Therefore, the ring coil, which changes its resonant frequency with the variable capacitor, can send its mouse sensor information to the wristband coil in the frequency-keying way. Moreover, the semi-passive signal transmission of the ring coil, which modifies the inductive field from the wristband coil in the load-modulation way, can save the power for signal transmission unlike the active signal transmission.
@@ 3.2. Ring
The ring coil consists of a -turned resonant coil with distributed capacitors and the mouse module equipped with a small trackball (EVQWJN007, Panasonic) supporting up, down, right, left scrolling and pressing interactions. When users scroll or press the trackball in the mouse module, the trackball either i) converts the scroll rotation of the trackball into rotation of cylindrical magnets with alternating polarity distribution or ii) turns on the tactile switch reacting to the user press of the trackball (see Figure 3b). The magnetic switch (CT8132BL, Allegro MicroSystems) below each magnet sends the magnet rotation data to an ultra-low-power MCU (STM32L011F4U6, STM) along with the ON/OFF data of the tactile switch, and then, according to the input type, the MCU applies a different corresponding voltage to a varactor diode (SMV1253, Skyworks) by a digital-potentiometer-based voltage divider circuit (AD5160, Analog Devices). Note that scrolling diagonally results in the simultaneous scrolling of adjacent horizontal and vertical magnets. For instance, when scrolling diagonally to the upper right, both the up and right magnet move concurrently. The varactor diode, which varies its capacitance from to based on the applied voltage, can change the by being connected in parallel to one of the distributed capacitors in the ring coil. In total, the ring coil provides five physical inputs: none, scroll down, up, left, right, and press. The resonant frequency for each input is set to ,, , , and , respectively. The inductance () and resistance () of the ring coil at are and , respectively. Note that the ring coil is tilted with approximately , mitigating the misalignment between the ring and wristband coils when the hand is grasped for thumb-to-index input.
@@ 3.3. Wristband
The wristband coil consists of a -turned flexible resonant coil mounted on a 3D-printed flexible wristband and the readout board connected to a mobile vector network analyzer (VNA) (NanoVNA-H4). To recognize the ring’s mouse state, the wristband coil detects the ring’s frequency peak () by monitoring the frequency sweep signal ranging from to with an input power of , the frequency bandwidth of , the sampling number of , and the sampling rate of fps. The wristband created by FDM 3D-printing of TPU filament is designed in the ellipse shape with width, height, and length to fit in the middle-sized adult hands. The wristband coil is connected to the readout board via two pin magnetic connectors (DIY Magnetic Connector, Adfruit). The reader board consists of the bridge circuit composed of six amplifiers (LTC6228, Analog Devices), magnetic connectors, and an external PC (Macbook Air). The NanoVNA sends the power spectrum of the frequency response to the connected external PC, which analyzes the frequency peak in the power spectrum, similar to the same peak detection algorithm as picoRing (Takahashi et al., 2024). In total, the resonant frequency, inductance () and resistance () of the wristband coil connected with eighteen distributed capacitors and a resistor in series are , , and , respectively, and the measured power consumption of the bridge circuit and NanoVNA are () and (), respectively. The current wristband coil created by the flexible PCBs and TPU-based band would have the challenge of wearability and stretchability. The use of stretchable textile coil (Kanada et al., 2025; Takahashi et al., 2021) is promising for high biocompatibility for the skin.
@@ 4.1. Operation Time of picoRing mouse
1: We measured the actual battery lifespan for this case.
Next, we measured how long picoRing mouse can operate. To calculate the expected battery life during usage, we first connected the battery terminals of picoRing mouse to a digital multimeter (34410A, Keysight). The measured power consumption at during ACTIVE and STANDBY mode is measured to be approximately () and (), respectively. Then, we estimated the battery lifespan hours for three types of different ACTIVE operation time of picoRing mouse. For the reference, we measured the battery lifespan hours when picoRing mouse operates continuously in ACTIVE mode. The battery lifespan is calculated from when the LiPo battery is fully charged at until it drops below (see Figure 5). To maintain the ACTIVE mode, the MCU is instructed to keep the digital voltage divider ON, i.e., shifting up the resonant frequency of the ring-shaped coil from (i.e., the digital voltage divider is OFF) to . Note that the magnetic scroll detector is also kept running continuously. The measured time is approximately hours with the curved Lipo battery (UFX150732).
Based on the result of power consumption and battery lifespan in continuous ACTIVE mode, Table 3 shows the expected battery lifespan for the different operation hours (i.e., 4, 8, or 24 hours per day). The lifespan hour is estimated based on two types of the battery capacity: 20 mAh and 27 mAh, both of which is widely used in commercially-available smart ring products. Based on the practical usage time of a mouse without causing tendinitis, which is approximately 4-8 hours per day (Mikkelsen et al., 2007; Gerr et al., 2002; Wolff et al., 2018), picoRing mouse has the potential to operate for approximately (8 hrs/day)- (4 hrs/day) hours, or over a few weeks or one month, on a single charge of mAh battery.

### v2

# Ultra-low-power ring-based wireless tinymouse

@@ Abstract.
The image shows the picoRing mouse, an ultra-low-power wireless ring mouse. On the left, a close-up of the ring mouse is displayed, highlighting its compact design with electronic components. In the center, a person is wearing the ring mouse on their finger, with an additional electronic module on their wrist. On the right, a digital interface shows a blue ring on a screen, indicating interaction with the device. Text on the image reads ”picoRing mouse, 30-500 uW, 5 gram.”
@@ 1. INTRODUCTION
Mouse devices are basic tools for simple, fast pointing with computers (Villar et al., 2009). Integrating tinymouse functionality into wearable form factors such as wristbands, glasses and rings (e.g., Apple Watch, Galaxy Ring) promises always-available essential interactions with wearable computing platforms (Xu et al., 2023; Kienzle et al., 2021). Especially, the wireless ring-formed input devices worn on the index finger offer reliable detection of even subtle, privacy-preserved thumb-to-index finger inputs, unlike the wristband and glasses (Takahashi et al., 2024; Kim et al., 2024). However, the physical constraint of the tiny ring structure requires the equipment of tens of small batteries, challenging the continuous operation of tens of -class power-consuming wireless communication modules. For example, prior wireless rings support below 1-10 hours of continuous wireless communication with Bluetooth Low Energy (BLE), limiting the ring’s communication usage to intermittent data transmission (Zhou et al., 2023; Kim et al., 2024; Shen et al., 2024; Jing et al., 2012). Such an operation is suitable for periodic healthcare monitoring around the rings (e.g., Oura Ring, Galaxy Ring), but not appropriate for the ring-based input interface that requires long-term, real-time communication with other wearables (e.g., smartwatch, HMD).
This paper presents picoRing mouse, enabling an ultra-low-power ring-shaped wireless tinymouse with a ring-to-wristband coil-based impedance sensing (see Figure 1). picoRing mouse is inspired by a coil-based sensitive impedance sensing named passive inductive telemetry (PIT) (Takahashi et al., 2024, 2020). Unlike long-range electromagnetic communication such as BLE and RF backscatter, PIT constructs a short-range but ultra-low-powered inductive link between a pair of a ring coil and a wristband coil. Since the ring coil can send its sensor information to the wristband coil by simply modifying the inductive field generated from the wristband coil, the ring coil does not need active signal transmission with power-hungry communication modules. With the combination of PIT with the ring-based mouse module, picoRing mouse enables the ultra-low-powered tinymouse ring with maximum power consumption of approximately , enabling continuous operation of approximately (24hrs use/day)- (4hrs use/day) days on a single charge of a curved Lipo battery. Furthermore, the shift from a previous full-passive ring design in original picoRing (Takahashi et al., 2024) to our semi-passive ring design by integrating an ultra-low-power MCU enables accurate detection of subtle mouse inputs without any physical noise issues (e.g. chattering) and demonstrates picoRing’s practical use in real-world scenarios. To facilitate replication and provide comprehensive implementation details, we have open-sourced the current picoRing mouse: https://github.com/KawaharaLab/picoRing_mouse.
@@ 2. RELATED WORK
The schematic illustrates the passive inductive telemetry (PIT) system for a ring-to-wristband wireless connection in the picoRing device. At the top, a diagram shows a hand wearing a ring coil and a wristband coil connected to a readout board. Green arrows represent the inductive field between the ring and wristband. Below, circuit diagrams depict the ring coil’s on/off states and its interaction with the wristband coil. A graph shows frequency shifts corresponding to the coil states. Annotations explain the role of distributed capacitors in increasing inductance at high frequencies.
@@ 2.1. Wearable Mouse
The table compares technical specifications of various wireless ring-based input devices. Columns include the ring name, components, wireless connection type, ring’s maximum power consumption in mW, available input interfaces, and user/session independence. The picoRing mouse is highlighted as the first device achieving ultra-low power consumption below 1 mW, specifically 0.45 mW, using a semi-PIT connection. Other devices listed include ElectroRing, Ring-a-Pose, OmniRing, and more, with varying power consumption and input methods.
@@ 2.2. Wireless Ring
Ring-formed mouse devices have long been investigated within the HCI community. The exploration of the previous mouse ring can be divided into two categories: i) ring-based sensing techniques such as computer vision (Sun et al., 2021; Chan et al., 2015; Kienzle and Hinckley, 2014), inertial sensors (Shen et al., 2024; Liang et al., 2021), ultrasound (Yu et al., 2024; Zhang et al., 2017), and electromagnetic wave (Chen et al., 2022; Waghmare et al., 2023) and ii) low-powered wireless ring design using RF backscatter (Liu et al., 2013; Ensworth and Reynolds, 2017; Bainbridge and Paradiso, 2011), near-field communication (NFC) (Zhan et al., 2024; Lee et al., 2022; Liang and Guo, 2021), intrabody communication (Fukumoto and Tonomura, 1997; Varga et al., 2018), and magnet-based tracking (Ashbrook et al., 2011). The ring-based finger sensing demonstrates high-fidelity microgesture recognition compared to the wristband- and glasses-based finger sensing (Kienzle et al., 2021). However, the long-term operation of the low-powered wireless rings is still challenging (Wagih et al., 2023). Specifically, -class RF backscatter and magnet-mounted ring need a large antenna and a bulky magnet, which is hard to fit in the small ring (Naderiparizi et al., 2018; Ashbrook et al., 2011). Moreover, the magnet-based tracking via exposed static magnetic field is easily influenced by the surrounding magnetic noises such as electrical appliances, impairing the tracking accuracy of the magnet position or orientation (Wang et al., 2022). -class NFC supports small tag design, but the communication range is a few centimeter-scale (Gummeson et al., 2014; Takahashi et al., 2025; Zhang et al., 2022); such a short-range link cannot construct ring-to-wristband communication. Intrabody communication (Fukumoto and Tonomura, 1997; Varga et al., 2018), which uses capacitive body as a signal path, can transmit signal with -class ultra-low-power throughout the body (Li et al., 2021), but both the RF presence of other electronic devices and the daily items such as wooden desk and wall can impair the capacitive signal path, causing signal degradation, reduced transmission range, or even data loss (Kienzle et al., 2021).
@@ 2.3. Passive Inductive Telemetry
The design overview of the picoRing mouse is shown in three parts. (a) A schematic depicts the ring coil and wristband coil connected to a readout board, illustrating the inductive field and resonant frequency shift. (b) Illustrations show the mouse sensor module with magnetic and tactile switches, connected to a microcontroller (MCU). The diagram explains how input is converted into electrical signals, highlighting active and standby current states. (c) The circuit diagram displays components like the ring coil, wristband coil, and readout board, detailing the vector network analysis process for frequency measurement.
@@ 3. SYSTEM DESIGN
picoRing mouse consists of two main components (see Figure 3a): 1) a ring coil with a mouse module, which changes its frequency response based on the thumb-to-index inputs, and 2) a wristband coil that detects the unique peak in the frequency response generated by the ring’s resonance. The working principle of picoRing mouse is as follows: First, the wristband coil generates a weak inductive field to couple with the nearby ring coil. When the user scrolls or presses the mouse module mounted on the ring, the MCU connected to the mouse module detects either the scrolling or pressing via the magnet or tactile switch, transforming the mouse input into a unique frequency response (i.e., the change in the ring’s resonant frequency) via a voltage-controlled variable capacitor (see Figure 3b). Owing to the inductive coupling, the wristband coil connected to a sensitive impedance measurement module obtains the frequency response. Since the ring’s frequency response appears as a sharp frequency peak at the ring’s resonant frequency in the frequency characteristics of the wristband’s impedance, the wristband coil enables to recognize the mouse input (i.e., pressing and 2D scrolling) as a unique peak output.
@@ 3.1. Ring-to-wristband Wireless Communication
picoRing mouse extends the PIT system in a semi-passive manner. While the prior PIT features the full-passive design of the ring coil by modifying the ring-to-wristband inductive field with a full-passive modulation circuit composed of tactile switches. In contrast, picoRing mouse employs the semi-passive modulation approach, which modifies the inductive field with a low-powered digital modulation circuit composed of a battery-assisted voltage-controlled variable capacitor. Note that the semi-passive modulation approach does not use the active signal transmission circuit. Unlike full-passive modulation, the semi-passive modulation can transmits the sensor information as digital signals. Specifically, picoRing mouse uses a simple frequency-shift keying, which encodes the multi-modal mouse input into a corresponding frequency shift based on the voltage-controlled variable capacitor.
Let us explain how the ring coil can influence the wristband coil in the frequency-shifting manner. When the ring coil resonates at inductively couples with the wristband coil, the input impedance of the wristband coil, , can be expressed as follows:
where and are the total impedance of the wristband and ring coils, respectively, is the angular frequency, is the mutual inductance between the ring and wristband coils. can be simplified as at because is in the resonance state. As a result, Equation 2 can be also simplified at as follows:
Equation 5 indicates increases significantly around , resulting in the sharp peak in the frequency characteristics of around . However, similar to picoRing (Takahashi et al., 2024), the extremely weak inductive coupling makes too small to detect through a standard impedance measurement circuit.
@@
where is the input voltage of the bridge circuit, is the amplifier factor. Note that is valid under condition because has the large imaginary part at non-resonant frequency and is so small in the situation where the ring coil is away from the wristband coil. Equation 8 indicates that the wristband coil can reliably detect the ring’s resonant frequency through output () of the bridge circuit. Therefore, the ring coil, which changes its resonant frequency with the variable capacitor, can send its mouse sensor information to the wristband coil in the frequency-keying way. Moreover, the semi-passive signal transmission of the ring coil, which modifies the inductive field from the wristband coil in the load-modulation way, can save power for signal transmission unlike the active signal transmission.
@@ 3.2. Ring
The photograph shows the prototype of the picoRing mouse. (a) Displays the ring coil on a 45 × 14 mm flexible PCB with labeled components: magnetic mouse, mouse sensor module, MCU, varactor diode, digital potentiometer, and switching regulator. Below, images show the ring coil in straight and angled types, mounted on a 3D-printed structure. (b) Shows the wristband coil on flexible PCBs connected to a readout board, including a balanced bridge circuit, NanoVNA, and external PC. Components like distributed capacitors and power supply are labeled.
The graph shows battery voltage (V) over time (hours) for the wireless ring. The blue line represents battery voltage decreasing from 4.2V to just below 3.7V over 192 hours. A horizontal orange line at 3.7V indicates a reference voltage level. The plot illustrates a gradual decline in voltage over time, with a sharper drop near the end.
The ring coil consists of an -turned resonant coil with distributed capacitors and the mouse module equipped with a small trackball (EVQWJN007, Panasonic) supporting up, down, right, left scrolling and pressing interactions. When users scroll or press the trackball in the mouse module, the trackball either i) converts the scroll rotation of the trackball into rotation of cylindrical magnets with alternating polarity distribution or ii) turns on the tactile switch reacting to the user press of the trackball (see Figure 3b). The magnetic switch (CT8132BL, Allegro MicroSystems) below each magnet sends the magnet rotation data to an ultra-low-power MCU (STM32L011F4U6, STM) along with the ON/OFF data of the tactile switch, and then, according to the input type, the MCU applies a different corresponding voltage to a varactor diode (SMV1253, Skyworks) by a digital-potentiometer-based voltage divider circuit (AD5160, Analog Devices). The varactor diode, which varies its capacitance from to based on the applied voltage, can change the by being connected in parallel to one of the distributed capacitors in the ring coil.
The ring coil provides five physical inputs: none, scroll down, up, left, right, and press. The resonant frequency for each input is set to , , , , , and , respectively, within the impedance-balanced frequency band. The inductance () and resistance () of the ring coil at are and , respectively. Note that the ring coil is tilted with approximately , mitigating the misalignment between the ring and wristband coils when the hand is grasped for thumb-to-index input. Scrolling diagonally results in the simultaneous scrolling of adjacent horizontal and vertical magnets. For instance, when scrolling diagonally to the upper right, both the up and right magnets move concurrently. The MCU decodes this diagonal scrolling into vertical and horizontal scrolling in turn.
@@ 3.3. Wristband
The wristband coil consists of a -turned flexible resonant coil mounted on a 3D-printed flexible wristband and the readout board connected to a mobile vector network analyzer (VNA) (NanoVNA-H4). To recognize the ring’s mouse state, the wristband coil detects the ring’s frequency peak () by monitoring the frequency sweep signal ranging from to with an input power of , the frequency bandwidth of , the sampling number of , and the sampling rate of fps. This setup for fast VNA readout enables the reliable detection of subtle and fast finger scrolling but, the frequency peak sometimes shifts because the large frequency step resolution of the NanoVNA without noise reduction sometimes fluctuates the frequency peak data. To address this, we’ve designed the ring’s six resonant frequencies with a enough frequency space of , as described in § 3.2.
The wristband created by FDM 3D-printing of TPU filament is designed in the ellipse shape with width, height, and length to fit in the middle-sized adult hands. The wristband coil is connected to the readout board via two pin magnetic connectors (DIY Magnetic Connector, Adfruit). The reader board consists of the bridge circuit composed of six amplifiers (LTC6228, Analog Devices), magnetic connectors, and an external PC (Macbook Air). The NanoVNA sends the power spectrum of the frequency response to the connected external PC, which analyzes the frequency peak in the power spectrum, similar to the same peak detection algorithm as picoRing (Takahashi et al., 2024). In total, the resonant frequency, inductance () and resistance () of the wristband coil connected with eighteen distributed capacitors and a resistor in series are , , and , respectively, and the measured power consumption of the bridge circuit and NanoVNA are () and (), respectively. The current wristband coil created by the flexible PCBs and TPU-based band would have the challenge of wearability and stretchability. The use of stretchable textile coil (Kanada et al., 2025; Takahashi et al., 2021) is promising for high biocompatibility for the skin.
@@ 4. TECHNICAL EVALUATION
The table lists power consumption of ring coil hardware components in standby and active modes, measured in microwatts (µW). Components include the MCU (STM32L011F4U6), mouse sensor module (4 × CT8132), digital voltage divider (MAX5394), and power management (TPS62843). Standby consumption ranges from 0.63 to 7.9 µW, while active consumption ranges from 0.63 to 252 µW. Estimated total is over 12 µW in standby and over 296 µW in active. Measured total at 4.2V is 25 µW in standby and 449 µW in active.
[…]
4.1. Operation Time of picoRing mouse
First, we measured how long picoRing mouse can operate. To calculate the expected battery life during usage, we first connected the battery terminals of picoRing mouse to a digital multimeter (34410A, Keysight). The measured power consumption at during ACTIVE and STANDBY mode is measured to be approximately () and (), respectively. Then, we estimated the battery lifespan hours for three types of different ACTIVE operation time of picoRing mouse. For reference, we measured the battery lifespan hours when picoRing mouse operates continuously in ACTIVE mode. The battery lifespan is calculated from when the LiPo battery is fully charged at until it drops below (see Figure 5). To maintain the ACTIVE mode, the MCU is instructed to keep the digital voltage divider ON, i.e., shifting up the resonant frequency of the ring-shaped coil from (i.e., the digital voltage divider is OFF) to . Note that the magnetic scroll detector is also kept running continuously. The measured time is approximately hours, i.e., 7 days with the fully-charged Lipo battery (UFX150732).
Based on the result of power consumption and battery lifespan in continuous ACTIVE mode, Table 3 shows the expected battery lifespan for the different operation hours (i.e., 4, 8, or 24 hours per day). The lifespan hour is estimated based on two types of battery capacity: 20 mAh and 27 mAh, both of which is widely used in commercially-available smart ring products. Based on the practical usage time of a mouse without causing tendinitis, which is approximately 4-8 hours per day (Mikkelsen et al., 2007; Gerr et al., 2002; Wolff et al., 2018), picoRing mouse has the potential to operate for approximately (8 hrs/day)- (4 hrs/day) hours, or over a few weeks or one month, on a single charge of mAh battery.
Figure 6 showcases the Signal-to-Noise Ratio (SNR) of the picoRing mouse in controlled environments. (a) A time-series graph illustrates the SNR for scrolling and pressing interactions among three users, with frequency on the y-axis and SNR values annotated for each interaction type: no input, scroll down, scroll up, scroll left, scroll right, and press. (b) Two images show different angles and distances between the ring and wristband coils, accompanied by a graph plotting SNR against distance at various angles (-30°, 0°, 30°, 30° no tilt, and 60°). (c) A graph displays SNR as a function of input power to the wristband coil, with input power on the x-axis and SNR on the y-axis, highlighting a peak SNR at a specific input power level.
Figure 7 displays the Signal-to-Noise Ratio (SNR) of the picoRing mouse during three electromagnetically noisy scenarios in real-world wearable computing. The top row shows images of each scenario: a person sitting on public transport, a group setting with multiple people using electronic devices, and a person driving a car while using the device. The bottom row contains corresponding SNR graphs for each scenario, with frequency on the y-axis and annotated SNR values of 27.3±2.1, 25.7±0.5, and 27.0±2.8. A color bar on the right indicates the SNR scale from 10 to 30.
Figure 8 illustrates application examples of the picoRing mouse. The images depict three scenarios: on an airplane, a person using the device to interact with a projected display; in a cityscape, a user interacting with a virtual map and keyboard; and on public transport, a person using the device to read a virtual document. Each scenario includes a close-up view of the picoRing mouse on the user’s finger, highlighting its integration with wearable technology.